OpenStudy (anonymous):

Find all the angles between 0 degree and 360 degree which satisfy 4sin x = 2 cos2x - 3

OpenStudy (anonymous):

Do i need to use double angle formula?

OpenStudy (anonymous):

which should i use?

OpenStudy (anonymous):

note that \(\cos 2x=1-2 \sin^2 x\)

OpenStudy (anonymous):

noted thanks.

OpenStudy (anonymous):

very well...:)

OpenStudy (anonymous):

carry on please /

OpenStudy (anonymous):

well make a quadratic equation in terms of \(\sin x\)

OpenStudy (anonymous):

i get this 2 sin^2 x + 4 sin x = -1

OpenStudy (anonymous):

then i take out common factor is that right?

OpenStudy (anonymous):

emm...yeah solve it by quadratic formula or factoring...which one u are comfortable with

OpenStudy (anonymous):

My answer are x = 270 degree and 315 degree

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

let me check it...

OpenStudy (anonymous):

\(4\sin x = 2 \cos 2x - 3 \) -----> \(4\sin x+4 \sin^2 x = -1\)

OpenStudy (anonymous):

or \( (2 \sin x+1)^2=0 \) and it gives \(\sin x=-1/2 \)

OpenStudy (anonymous):

so 210 and 350

OpenStudy (anonymous):

oh i see what you did there >.>

OpenStudy (anonymous):

can you wait for awhile coz i have another question :(

OpenStudy (anonymous):

i think u made a little mistake in making quadratic equation...